Mr. Anders and Ms. Rich each drove home from a business meeting. Mr. Anders traveled east at 100 kilometers per hour and Ms. Rich traveled west at 80 kilometers per house. In how many hours were they 100 kilometers apart?
Answer by mananth(16949)

they are going away from each other
speed of DISABLED_event_one= 100 mph
speed of other = 80 km/h
..
180 km/h is the effective speed.
distance = 100 km
time = d/r
t=100/180
t=5/9 hours
...
m.ananth@hotmail.ca